OpenStudy (anonymous):

Find the tr(A) such that (A^t- 5I)= B where I is a 2 x 2 identity matrix where B is 0 1 -1 0

OpenStudy (rational):

A^t - 5I = B A^t = B+5I

OpenStudy (rational):

find A^t first

OpenStudy (freckles):

Well the tr(A)=tr(the transpose of A)

OpenStudy (freckles):

so just add B+5I and find the tr(transpose of A) and you are done

OpenStudy (anonymous):

well I didn't know I could do "algebra" with matrices. I thought normal rule didn't apply

OpenStudy (anonymous):

oh sorry! its (a^t-I5)^-1 :/